# Global Self-Attention with Exact Fourier Propagation for Phase-Only Far-Field Holography

### v1

@@ Abstract
Phase-only computer-generated holography (CGH) seeks a phase pattern for a spatial light modulator (SLM) whose propagated optical field reproduces a desired intensity distribution. In the far-field (Fraunhofer) regime, optical propagation reduces to a Fourier transform, such that each hologram pixel contributes to the entire reconstructed intensity distribution. When restricted to phase-only modulation, intensity must be shaped through global phase interference effects, making the inverse mapping from target intensity to phase highly non-linear and sensitive to local minima. We present a proof-of-concept physics-in-the-loop approach in which a transformer maps a target intensity image to a phase-only SLM field and is trained end-to-end through exact FFT-based propagation embedded directly within optimization. We further observe that patch tokenization strongly shapes the optimization geometry: coarse tokenization acts as an implicit spectral regularizer that stabilizes training and suppresses checkerboard-like attractors, while finer tokenization increases spatial degrees of freedom but benefits from curriculum or hierarchical refinement. Despite training on limited primitives and restricted digit subsets, the learned generator exhibits out-of-distribution (OOD) generalization to unseen digits and hand-drawn target patterns. These results suggest that transformer architectures, whose self-attention enables global token interactions, are a natural fit for far-field holography and provide a viable foundation for scalable physics-grounded hologram generation.
@@ 2.5 Training Data and Curriculum
To encourage broad generalization beyond any single dataset, we train on a mixed synthetic + real-intensity curriculum that incrementally increases target complexity. The final stage training set aggregates multiple categories including procedurally generated patterns (e.g., structured basis-like stimuli, symmetric dot patterns, sparse and dense random patterns, and gratings), as well as large sets of synthetic shapes, doodles, and band-limited smooth textures. Additionally, digit-like and apparel-like targets are included via MNIST and Fashion-MNIST samples.[MNISTweb][Xiao2017FashionMNIST] In the final stage, the training set composition is presented in Table 1.
@@ 3.1 Quantitative Results
To probe OOD behavior on digit targets, we additionally sample $n=10$ examples per digit from the full MNIST test set (digits 099) and report mean MSE per digit. Despite training that included a digit-restricted curriculum tier (and later broad synthetic targets), the trained model produces recognizable far-field digit structure across all classes. The per-digit OOD mean MSE is presented in Table 3. We observe the lowest OOD error on digit 6 in this sampled evaluation, consistent with the curriculum including a digit-specific tier.
@@ 3.3 Generalization to Custom Hand-Drawn Targets
(a) Example 1 (MSE = 1.0848)
(b) Example 2 (MSE = 1.4815)
(c) Example 3 (MSE = 1.1624)
(d) Example 4 (MSE = 1.7201)
@@ 4 Discussion and Conclusion
The results demonstrate that transformer-based architectures trained with explicit physics-in-the-loop optimization can stably generate phase-only far-field holograms under exact FFT propagation. Across synthetic patterns, digit subsets, and manually constructed targets, the model converges without collapse and produces coherent reconstructions that preserve global structure. The observed OOD behavior further suggests the network learns aspects of the inverse diffraction mapping rather than merely memorizing training exemplars.

### v2

@@ Abstract
Phase-only computer-generated holography (CGH) seeks a phase pattern for a spatial light modulator (SLM) whose propagated optical field reproduces a desired intensity distribution. In the far-field (Fraunhofer) regime, optical propagation reduces to a Fourier transform, such that each hologram pixel contributes to the entire reconstructed intensity distribution. When restricted to phase-only modulation, intensity must be shaped through global phase interference effects, making the inverse mapping from target intensity to phase highly non-linear and sensitive to local minima. We present a proof-of-concept physics-in-the-loop approach in which a transformer maps a target intensity image to a phase-only SLM field and is trained end-to-end through exact FFT-based propagation embedded directly within optimization. We further observe that patch tokenization strongly shapes the optimization geometry: coarse tokenization acts as an implicit spectral regularizer that stabilizes training and suppresses checkerboard-like attractors, while finer tokenization increases spatial degrees of freedom but benefits from curriculum or hierarchical refinement. Despite training on limited primitives and a single digit class (only digit 6), the learned generator exhibits out-of-distribution (OOD) generalization to unseen digits and hand-drawn target patterns. These results suggest that transformer architectures, whose self-attention enables global token interactions, are a natural fit for far-field holography and provide a viable foundation for scalable physics-grounded hologram generation.
@@ 2.5 Training Data and Curriculum
To encourage broad generalization beyond any single dataset, we train on a mixed synthetic + real-intensity curriculum that incrementally increases target complexity. The final stage training set aggregates multiple categories including procedurally generated patterns (e.g., structured basis-like stimuli, symmetric dot patterns, sparse and dense random patterns, and gratings), as well as large sets of synthetic shapes, doodles, and band-limited smooth textures. Additionally, digit-like and apparel-like targets are included via MNIST and Fashion-MNIST samples.[MNISTweb][Xiao2017FashionMNIST] For the digit-specific curriculum stage, training was restricted to samples of a single digit class (digit 6) from MNIST. No other digit classes were included during training. This design allows explicit evaluation of cross-class generalization to unseen digits. In the final stage, the training set composition is presented in Table 1.
@@ 3.1 Quantitative Results
To probe out-of-distribution behavior on digit targets, we evaluate the model on all ten MNIST digit classes (0–9), despite the training curriculum including only a single digit class (digit 6). The model therefore has no exposure during training to the remaining nine digit categories. We sample $n=10$ examples per digit from the MNIST test set and report mean MSE per class. The trained model produces recognizable far-field digit structure across all classes. The per-digit OOD mean MSE is presented in Table 3. We observe the lowest OOD error on digit 6 in this sampled evaluation, consistent with the curriculum including a digit-specific tier.
@@ 3.3 Generalization to Custom Hand-Drawn Targets
(a) Example 1
(b) Example 2
(c) Example 3
(d) Example 4
@@ 4 Discussion and Conclusion
The results demonstrate that transformer-based architectures trained with explicit physics-in-the-loop optimization can stably generate phase-only far-field holograms under exact FFT propagation. Across synthetic patterns, MNIST digit 6 subset, and manually constructed targets, the model converges without collapse and produces coherent reconstructions that preserve global structure. The observed OOD behavior further suggests the network learns aspects of the inverse diffraction mapping rather than merely memorizing training exemplars.